# A Variational Principle for Extended Irreversible Thermodynamics: Heat Conducting Viscous Fluids

### v1

@@ 3.3 Variational formulation in the spatial description
Consider a viscous heat conducting fluid with nonequilibrium Lagrangian density $\widehat{\mathfrak{L}}(\boldsymbol{\varphi},\dot{\boldsymbol{\varphi}},\nabla{\boldsymbol{\varphi}},\varrho,S,\boldsymbol{S},\boldsymbol{Q})$ expressed in terms of the configuration map $\boldsymbol{\varphi}$ and its first derivatives, in terms of the mass and entropy densities $\varrho$ and $S$, and in terms of the thermodynamic fluxes $\boldsymbol{S}$ and $\boldsymbol{Q}$.
@@
    $\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\frac{d}{dt}\frac{\partial\widehat{\mathfrak{L}}}{\partial\dot{\boldsymbol{\varphi}}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{\varphi}}=\operatorname{DIV}\Big{(}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\nabla{\boldsymbol{\varphi}}}+\boldsymbol{P}\Big{)}\\ \displaystyle-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}(\dot{S}+\operatorname{DIV}\boldsymbol{J})-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}=\boldsymbol{S}:\boldsymbol{D}(\dot{\boldsymbol{\varphi}})+\boldsymbol{J}\cdot\nabla\frac{\partial\widehat{\mathfrak{L}}}{\partial S},\end{array}\right.$
@@
    $\dot{\Gamma}=-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\qquad\text{and}\qquad\dot{\Sigma}=\dot{S}+\operatorname{DIV}\boldsymbol{J}.$
  • Spatial description: When the nonequilibrium Lagrangian density $\widehat{\mathfrak{L}}$ satisfies the material covariance assumption, it has an associated nonequilibrium Eulerian Lagrangian density ^$\widehat{\ell}\,(\boldsymbol{u},\rho,s,\boldsymbol{\sigma},\boldsymbol{q})$ and (3.1)–(3.3) yields the Eulerian variational formulation (3.28)–(3.30). This variational principle yields the equations of motion
    $\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\partial_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\widehat{\ell}}{\partial\rho}+s\nabla\frac{\partial\widehat{\ell}}{\partial s}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\nabla\boldsymbol{\sigma}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\nabla\boldsymbol{q}+\operatorname{div}(\boldsymbol{\sigma}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\tau}_{\boldsymbol{q}})\\ \vspace{0.2cm}\displaystyle-\frac{\partial\widehat{\ell}}{\partial s}(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla\frac{\partial\widehat{\ell}}{\partial s}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\\ \displaystyle\bar{D}_{t}\rho=0.\end{array}\right.$
    together with the insulated boundary condition $\boldsymbol{j}\cdot\boldsymbol{n}=0$. Furthermore, the variational principle also determines the following relationships for the variables $\gamma$ and $\varsigma$ in (3.28)–(3.3):
    $D_{t}\gamma=-\frac{\partial\widehat{\ell}}{\partial s}\qquad\text{and}\qquad\bar{D}_{t}\varsigma=\bar{D}_{t}s+\operatorname{div}\boldsymbol{j}.$
Equation for standard Lagrangians.
$\left\{\begin{array}[]{l}\vspace{0.2cm}\displaystyle\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p+\operatorname{div}(\boldsymbol{\tau}_{\boldsymbol{q}}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\sigma})\\ \vspace{0.2cm}\displaystyle T(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\\ \displaystyle\bar{D}_{t}\rho=0\end{array}\right.$ (3.42)
@@
$\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p+\operatorname{div}(\boldsymbol{\tau}_{\boldsymbol{b}}),\qquad\boldsymbol{\tau}_{\boldsymbol{b}}=\frac{\partial\varepsilon}{\partial\boldsymbol{b}}\cdot\boldsymbol{b}$
with the pressure given by $p=\rho\frac{\partial\varepsilon}{\partial\rho}+s\frac{\partial\varepsilon}{\partial s}-\varepsilon$. One notices the similarity between the mathematical form of $\boldsymbol{\tau}_{\boldsymbol{b}}$ and $\boldsymbol{\tau}_{\boldsymbol{q}}$ in terms of their respective energy density, with the slight difference between them due to the fact that $\boldsymbol{\sigma}$ is a tensorial density, whereas $\boldsymbol{b}$ is an ordinary tensor (if $\rho\boldsymbol{b}$ would be considered instead, then same mathematical form would be achieved). A comparable observation holds for $\boldsymbol{\tau}_{\boldsymbol{q}}$ in the case of a conservative continuum involving the tensorial transport of a vector field density. Despite these formal analogies, the physical nature of the stresses is fundamentally different. In the case of nonlinear elasticity, $\boldsymbol{\tau}_{\boldsymbol{b}}$ (at least in the equations above) arises from a reversible, conservative process, with $\boldsymbol{b}$ evolving according to a continuity-like equation. In contrast, the stresses $\boldsymbol{\tau}_{\boldsymbol{q}}$ and $\boldsymbol{\tau}_{\boldsymbol{\sigma}}$ in (3.43) describe extended irreversible processes, and the underlying fields do not satisfy such transport equations. This underscores the distinct physical origins of these stresses, despite their superficial mathematical resemblance. (3) It is possible to reorganise the terms in the momentum balance in (3.41) in a slightly different way as
@@ 3.4 Energy and entropy balance in the spatial description
$\widehat{e}=\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}\cdot\boldsymbol{u}-\widehat{\ell}.$
@@
$\frac{d}{dt}\int_{\mathcal{D}}s\,{\rm d}\mathbf{x}=\int_{\mathcal{D}}\frac{1}{T}\bigg{(}\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\bigg{)}{\rm d}\mathbf{x}-\int_{\partial\mathcal{D}}\boldsymbol{j}\cdot\boldsymbol{n}\,{\rm d}A.$
@@
$\bar{D}_{t}\varsigma=\frac{1}{T}\bigg{(}\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\bigg{)}.$
Balance laws for standard Lagrangians.
The energy and entropy balance laws for the system described by equation (3.42) take the following form
$\bar{D}_{t}\widehat{e}=\operatorname{div}\big{(}(-p\delta+\boldsymbol{\tau}_{\boldsymbol{q}}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\sigma})\cdot\boldsymbol{u}-\boldsymbol{j}T\big{)}$
$\bar{D}_{t}s+\operatorname{div}\boldsymbol{j}=\frac{1}{T}\Big{(}\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\Big{)}\geq 0$
@@
$D_{t}\widehat{e}=\frac{1}{2}D_{t}\rho|\boldsymbol{u}|^{2}+\rho\boldsymbol{u}\cdot D_{t}\boldsymbol{u}+\frac{\partial\varepsilon}{\partial\rho}D_{t}\rho+\frac{\partial\varepsilon}{\partial s}D_{t}s+\frac{\partial\varepsilon}{\partial\boldsymbol{q}}\cdot D_{t}\boldsymbol{q}$
@@
$D_{t}\widehat{e}=-\widehat{e}\operatorname{div}\boldsymbol{u}+\operatorname{div}(-p\boldsymbol{u}-T\boldsymbol{j})+\boldsymbol{u}\cdot\operatorname{div}(\boldsymbol{\tau}_{\boldsymbol{q}})-\frac{\partial\varepsilon}{\partial\boldsymbol{q}}\cdot(\mathfrak{D}_{t}\boldsymbol{q}-D_{t}\boldsymbol{q}).$
@@ 3.5 Phenomenology: Maxwell and Cattaneo-Christov laws
$\boldsymbol{\sigma}:\big{(}\operatorname{Def}\boldsymbol{u}-\beta\mathfrak{D}_{t}\boldsymbol{\sigma}\big{)}+\boldsymbol{q}\cdot\big{(}-\frac{1}{T}\nabla T-\alpha\mathfrak{D}_{t}\boldsymbol{q}\big{)}\geq 0,$
@@
Viscosity.
Comparison with the Newton-Stokes law yields from the first relation in (3.50) and particular choice of $\mathbb{L}$ an extended compressible Maxwell model with Truesdell rate
$\left\{\begin{array}[]{l}\displaystyle\vspace{0.2cm}\tau_{2}(\mathfrak{D}_{t}\boldsymbol{\sigma})^{(0)}+\boldsymbol{\sigma}^{(0)}=2\eta(\operatorname{Def}\boldsymbol{u})^{(0)}\\ \displaystyle\tau_{0}\frac{1}{3}\operatorname{Tr}\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{1}{3}\operatorname{Tr}\boldsymbol{\sigma}=\zeta\operatorname{div}\boldsymbol{u},\end{array}\right.$ (3.51)
where $\eta$ and $\zeta$ denote the shear and bulk viscosity coefficients, respectively, and the relaxation coefficients are given by $\tau_{2}=2\eta\beta$, $\tau_{0}=3\zeta\beta$. Here the superscript ${(0)}$ denotes the traceless part of the tensors. It is important to note that the time rate $\mathfrak{D}_{t}$ does not preserve the isotropic and traceless decomposition of symmetric tensor fields. As a result, system (3.51) exhibits coupling between the volumetric and deviatoric stresses. In more details, introducing the volumetric pressure component $p^{v}=\frac{1}{3}\operatorname{Tr}\boldsymbol{\sigma}$, equation (3.51) can be rewritten as
$\left\{\begin{array}[]{l}\displaystyle\vspace{0.2cm}\tau_{2}\left(\mathfrak{D}_{t}\boldsymbol{\sigma}^{(0)}+\frac{2}{3}\boldsymbol{\sigma}^{(0)}:(\operatorname{Def}\boldsymbol{u})^{(0)}\delta-2p^{v}(\operatorname{Def}\boldsymbol{u})^{(0)}\right)+\boldsymbol{\sigma}^{(0)}=2\eta(\operatorname{Def}\boldsymbol{u})^{(0)}\\ \displaystyle\tau_{0}\left(D_{t}p_{v}+\frac{1}{3}p^{v}\operatorname{div}\boldsymbol{u}-\frac{2}{3}\boldsymbol{\sigma}^{(0)}:(\operatorname{Def}\boldsymbol{u})^{(0)}\right)+p_{v}=\zeta\operatorname{div}\boldsymbol{u}.\end{array}\right.$ (3.52)
@@
Heat conduction.
Comparison of the second relation in (3.50) with Fourier’s law yields $\mathbb{K}=\kappa T\delta$, with $\kappa$ the thermal conductivity, so that this relation reduces to the Cattaneo-Christov heat flux model:
$\tau_{1}\mathfrak{D}_{t}\boldsymbol{q}+\boldsymbol{q}=-\kappa\nabla T,$

### v2

@@ 3.3 Variational formulation in the spatial description
Consider a viscous heat conducting fluid with nonequilibrium Lagrangian density
expressed in terms of the configuration map $\boldsymbol{\varphi}$ and its first derivatives, in terms of the mass and entropy densities $\varrho$ and $S$, and in terms of the thermodynamic fluxes $\boldsymbol{S}$ and $\boldsymbol{Q}$.
@@
    $\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\frac{d}{dt}\frac{\partial\widehat{\mathfrak{L}}}{\partial\dot{\boldsymbol{\varphi}}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{\varphi}}=\operatorname{DIV}\Big{(}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\nabla{\boldsymbol{\varphi}}}+\boldsymbol{P}\Big{)}\\ \displaystyle-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}(\dot{S}+\operatorname{DIV}\boldsymbol{J})-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{S}}:\dot{\boldsymbol{S}}-\frac{\partial\widehat{\mathfrak{L}}}{\partial\boldsymbol{Q}}\cdot\dot{\boldsymbol{Q}}=\boldsymbol{S}:\boldsymbol{D}(\dot{\boldsymbol{\varphi}})+\boldsymbol{J}\cdot\nabla\frac{\partial\widehat{\mathfrak{L}}}{\partial S},\end{array}\right$. (3.55)
@@
    $\dot{\Gamma}=-\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\qquad\text{and}\qquad\dot{\Sigma}=\dot{S}+\operatorname{DIV}\boldsymbol{J}.$ (3.56)
  • Spatial description: When the nonequilibrium Lagrangian density $\widehat{\mathfrak{L}}$ satisfies the material covariance assumption, it has an associated nonequilibrium Eulerian Lagrangian density $\widehat{\ell}\,(\boldsymbol{u},\rho,s,\boldsymbol{\sigma},\boldsymbol{q})$ and (3.2)–(3.4) yields the Eulerian variational formulation (3.40)–(3.42). This variational principle yields the equations of motion
    $\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\partial_{t}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}+\pounds_{\boldsymbol{u}}\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}=\rho\nabla\frac{\partial\widehat{\ell}}{\partial\rho}+s\nabla\frac{\partial\widehat{\ell}}{\partial s}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\nabla\boldsymbol{\sigma}-\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\nabla\boldsymbol{q}+\operatorname{div}(\boldsymbol{\sigma}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\tau}_{\boldsymbol{q}})\\ \vskip 5.69046pt\displaystyle-\frac{\partial\widehat{\ell}}{\partial s}(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}+\boldsymbol{j}\cdot\nabla\frac{\partial\widehat{\ell}}{\partial s}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\\ \displaystyle\bar{D}_{t}\rho=0,\end{array}\right$. (3.57)
    which incorporate the nonequilibrium stresses (3.51), together with the insulated boundary condition $\boldsymbol{j}\cdot\boldsymbol{n}=0$. Furthermore, the variational principle also determines the following relationships for the variables $\gamma$ and $\varsigma$ in (3.40)–(3.4):
    $D_{t}\gamma=-\frac{\partial\widehat{\ell}}{\partial s}\qquad\text{and}\qquad\bar{D}_{t}\varsigma=\bar{D}_{t}s+\operatorname{div}\boldsymbol{j}.$ (3.58)
Equation for standard Lagrangians. For the Lagrangian density of the standard form (3.25), one gets from (3.53) the system
$\left\{\begin{array}[]{l}\vskip 5.69046pt\displaystyle\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p+\operatorname{div}(\boldsymbol{\tau}_{\boldsymbol{q}}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\sigma})\\ \vskip 5.69046pt\displaystyle T(\bar{D}_{t}s+\operatorname{div}\boldsymbol{j})=\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\\ \displaystyle\bar{D}_{t}\rho=0\end{array}\right$. (3.59)
@@
$\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p+\operatorname{div}(\boldsymbol{\tau}_{\boldsymbol{b}}),\qquad\boldsymbol{\tau}_{\boldsymbol{b}}=2\frac{\partial\varepsilon}{\partial\boldsymbol{b}}\cdot\boldsymbol{b}$ (3.61)
with the pressure given by $p=\rho\frac{\partial\varepsilon}{\partial\rho}+s\frac{\partial\varepsilon}{\partial s}-\varepsilon$. One notices the similarity between the mathematical form of $\boldsymbol{\tau}_{\boldsymbol{b}}$ and $\boldsymbol{\tau}_{\boldsymbol{q}}$ in terms of their respective energy density, with the slight difference between them due to the fact that $\boldsymbol{\sigma}$ is a tensorial density, whereas $\boldsymbol{b}$ is an ordinary tensor (if $\rho\boldsymbol{b}$ would be considered instead, then the same mathematical form would be achieved). A comparable observation holds for $\boldsymbol{\tau}_{\boldsymbol{q}}$ in the case of a conservative continuum involving the tensorial transport of a vector field density. Despite these formal analogies, the physical nature of the stresses is fundamentally different. In the case of nonlinear elasticity, $\boldsymbol{\tau}_{\boldsymbol{b}}$ (at least in the equations above) arises from a reversible, conservative process, with $\boldsymbol{b}$ evolving according to a continuity-like equation. In contrast, the stresses $\boldsymbol{\tau}_{\boldsymbol{q}}$ and $\boldsymbol{\tau}_{\boldsymbol{\sigma}}$ in (3.60) describe extended irreversible processes, and the underlying fields do not satisfy such transport equations. This underscores the distinct physical origins of these stresses, despite their superficial mathematical resemblance. (3) It is possible to reorganise the terms in the momentum balance in (3.53) in a slightly different way as
@@
Remark 3.5 (Evolution equations for the fluxes).
In line with the philosophy of extended irreversible thermodynamics, the evolution equations for the fluxes $\boldsymbol{q}$ and $\boldsymbol{\sigma}$ are of phenomenological origin. They are constructed to fulfill the second law of thermodynamics within the extended framework. Accordingly, these evolution equations will be addressed later in §3.5 in the context of phenomenological modeling.
@@ 3.4 Energy and entropy balance in the spatial description
$\widehat{e}=\frac{\partial\widehat{\ell}}{\partial\boldsymbol{u}}\cdot\boldsymbol{u}-\widehat{\ell}$. (3.64)
@@
$\frac{d}{dt}\int_{\mathcal{D}}s\,{\rm d}\mathbf{x}=\int_{\mathcal{D}}\frac{1}{T}\bigg{(}\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\bigg{)}{\rm d}\mathbf{x}-\int_{\partial\mathcal{D}}\boldsymbol{j}\cdot\boldsymbol{n}\,{\rm d}A.$ (3.67)
@@
$\bar{D}_{t}\varsigma=\frac{1}{T}\bigg{(}\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{\partial\widehat{\ell}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\bigg{)}.$ (3.68)
Balance laws for standard Lagrangians. The energy and entropy balance laws for the system described by equation (3.59) take the following form
$\bar{D}_{t}\widehat{e}=\operatorname{div}\big{(}(-p\delta+\boldsymbol{\tau}_{\boldsymbol{q}}+\boldsymbol{\tau}_{\boldsymbol{\sigma}}+\boldsymbol{\sigma})\cdot\boldsymbol{u}-\boldsymbol{j}T\big{)}$ (3.69)
$\bar{D}_{t}s+\operatorname{div}\boldsymbol{j}=\frac{1}{T}\Big{(}\boldsymbol{\sigma}:\nabla\boldsymbol{u}-\boldsymbol{j}\cdot\nabla T-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{\sigma}}:\mathfrak{D}_{t}\boldsymbol{\sigma}-\frac{\partial\widehat{\varepsilon}}{\partial\boldsymbol{q}}\cdot\mathfrak{D}_{t}\boldsymbol{q}\Big{)}\geq 0$ (3.70)
@@
$D_{t}\widehat{e}=\frac{1}{2}D_{t}\rho|\boldsymbol{u}|^{2}+\rho\boldsymbol{u}\cdot D_{t}\boldsymbol{u}+\frac{\partial\varepsilon}{\partial\rho}D_{t}\rho+\frac{\partial\varepsilon}{\partial s}D_{t}s+\frac{\partial\varepsilon}{\partial\boldsymbol{q}}\cdot D_{t}\boldsymbol{q}$ (3.72)
@@
$D_{t}\widehat{e}=-\widehat{e}\operatorname{div}\boldsymbol{u}+\operatorname{div}(-p\boldsymbol{u}-T\boldsymbol{j})+\boldsymbol{u}\cdot\operatorname{div}(\boldsymbol{\tau}_{\boldsymbol{q}})-\frac{\partial\varepsilon}{\partial\boldsymbol{q}}\cdot(\mathfrak{D}_{t}\boldsymbol{q}-D_{t}\boldsymbol{q}).$ (3.73)
@@ 3.5 Phenomenology: Maxwell and Cattaneo-Christov laws
$\boldsymbol{\sigma}:\big{(}\operatorname{Def}\boldsymbol{u}-\beta\mathfrak{D}_{t}\boldsymbol{\sigma}\big{)}+\boldsymbol{q}\cdot\big{(}-\frac{1}{T}\nabla T-\alpha\mathfrak{D}_{t}\boldsymbol{q}\big{)}\geq 0,$ (3.75)
@@
Viscosity. Comparison with the Newton-Stokes law yields from the first relation in (3.76) and particular choice of $\mathbb{L}$ an extended compressible Maxwell model with Truesdell rate
$\left\{\begin{array}[]{l}\displaystyle\vskip 5.69046pt\tau_{2}(\mathfrak{D}_{t}\boldsymbol{\sigma})^{(0)}+\boldsymbol{\sigma}^{(0)}=2\eta(\operatorname{Def}\boldsymbol{u})^{(0)}\\ \displaystyle\tau_{0}\frac{1}{3}\operatorname{Tr}\mathfrak{D}_{t}\boldsymbol{\sigma}+\frac{1}{3}\operatorname{Tr}\boldsymbol{\sigma}=\zeta\operatorname{div}\boldsymbol{u},\end{array}\right$. (3.77)
where $\eta$ and $\zeta$ denote the shear and bulk viscosity coefficients, respectively, and the relaxation coefficients are given by $\tau_{2}=2\eta\beta$, $\tau_{0}=3\zeta\beta$. Here the superscript (0){(0)} denotes the traceless part of the tensors. It is important to note that the time rate $\mathfrak{D}_{t}$ does not preserve the isotropic and traceless decomposition of symmetric tensor fields. As a result, system (3.77) exhibits coupling between the volumetric and deviatoric stresses. In more details, introducing the volumetric pressure component $p^{v}=\frac{1}{3}\operatorname{Tr}\boldsymbol{\sigma}$, equation (3.77) can be rewritten as
$\left\{\begin{array}[]{l}\displaystyle\vskip 5.69046pt\tau_{2}\left(\mathfrak{D}_{t}\boldsymbol{\sigma}^{(0)}+\frac{2}{3}\boldsymbol{\sigma}^{(0)}:(\operatorname{Def}\boldsymbol{u})^{(0)}\delta-2p^{v}(\operatorname{Def}\boldsymbol{u})^{(0)}\right)+\boldsymbol{\sigma}^{(0)}=2\eta(\operatorname{Def}\boldsymbol{u})^{(0)}\\ \displaystyle\tau_{0}\left(D_{t}p_{v}+\frac{1}{3}p^{v}\operatorname{div}\boldsymbol{u}-\frac{2}{3}\boldsymbol{\sigma}^{(0)}:(\operatorname{Def}\boldsymbol{u})^{(0)}\right)+p_{v}=\zeta\operatorname{div}\boldsymbol{u}.\end{array}\right$. (3.78)
@@
Heat conduction. Comparison of the second relation in (3.76) with Fourier’s law yields $\mathbb{K}=\kappa T\delta$, with $\kappa$ the thermal conductivity, so that this relation reduces to the Cattaneo-Christov heat flux model:
$\tau_{1}\mathfrak{D}_{t}\boldsymbol{q}+\boldsymbol{q}=-\kappa\nabla T,$ (3.80)
@@
Extensions of classical heat transfer equations also arise from frameworks other than extended irreversible thermodynamics, such as the Green-Naghdi theories of nonclassical heat conduction based on the concept of thermal displacement [32]. More recently, a variational formulation in the context of hyperbolic heat equation was considered in [12], also relying on the thermal displacement. In that context, the evolution equation governs the gradient of the thermal displacement, which, unlike the evolution equation for the heat flux $\boldsymbol{q}$ considered in the extended irreversible thermodynamics framework, does not give rise to entropy production. Cattaneo-type heat conduction, although of a different nature than the formulations discussed here, also emerges from a Godunov-type structure, as shown in [57].
Remark 3.7 (Causality).
It will be important to analyze under which conditions the extended heat-conducting viscous fluid model developed in this paper - namely equations (3.59) and (3.60), together with the phenomenological expressions derived in §3.5 - is hyperbolic, as this would ensure the causal character of the system. Unlike in the simpler setting of pure heat conduction, where a damped Fourier law directly yields a hyperbolic temperature equation, the general case of heat-conducting fluids considered here involves a strong coupling between the temperature and heat flux evolution equations. This coupling originates from the structure of the entropy balance equation in extended irreversible thermodynamics (see, for instance, the second equation in (3.59)), which - when rewritten in terms of temperature - retains significant dependence on both the heat flux and its time derivative. Although the resulting temperature equation still resembles a hyperbolic equation, this entanglement with the flux dynamics complicates a straightforward interpretation of causality. The causal or non-causal character of the system may also depend on the specific form of the nonequilibrium internal energy. The closest work in the literature appears to be [1], which, however, demonstrates the non-hyperbolic nature of heat-conducting non-viscous fluid models that couple the standard mass, momentum, and energy balance equations with the Cattaneo-Christov fluid model. As discussed above, that model differs from ours and lacks thermodynamic consistency. We postpone the detailed analysis of hyperbolicity and causality of our model for future work. Alternatively, a numerical investigation of this property could be undertaken by leveraging the variational formulation developed here to construct thermodynamically consistent numerical schemes, as was done in [19] for the classical irreversible thermodynamics of heat-conducting fluids.